OpenStudy (anonymous):

How do you combine and solve for y' in the following: cos(xy) * (xy' + y) = 2x - y'

OpenStudy (anonymous):

Original problem: Find y' of sin(xy) = (x^2)-y

OpenStudy (dumbcow):

move the y' terms to one side, then factor out the y'

OpenStudy (dumbcow):

\[y' (x \cos(xy) +1) = 2x-y \cos(xy)\]